OpenStudy (anonymous):

Maureen cut a 20 cm ribbon into exactly three pieces. The first piece is 3 cm shorter than the second piece. The third piece is 4 cm shorter than the second piece. Find the length of the shortest piece

OpenStudy (campbell_st):

let x be the 1st piece x - 3 is the 2nd piece (x -3) - 4 or x - 7 is the 3rd piece so 20 = x + x - 3 + x - 7 or 20 = 3x - 10 now just solve for x, to find the 1st piece...and then the other 2 hope it helps